OpenStudy (anonymous):

Segments with the lengths of 6, 8, and 10 units will form what type of triangle?

OpenStudy (anonymous):

3-4-5 right angled triangle

OpenStudy (phi):

with these type problems, use the pythagorean theorem a^2 + b^2 = c^2 to see if you have a right triangle

Parth (parthkohli):

This question is ambiguous IMHO. It can be a scalene triangle as well!

OpenStudy (phi):

it is easier if you memorize that legs of length 3,4 and 5 is a right triangle, and even better if you multiply each leg by the same number, it is still a right triangle. for example, if we multiply each leg by 2 we get 6,8,10 (which matches your problem) yes to Parth, it is a scalene, right triangle

Parth (parthkohli):

Yup. A triangle with sides \(a,b,c\) where \(c\) is the largest side will be a right triangle if \((ka)^2 + (kb)^2 = (kc)^2\)